OpenStudy (anonymous):

subtract. simplify, if possible. x - 2 --- --- x^2+5x+6 x^2+3x+2

myininaya (myininaya):

\[\frac{x}{(x+3)(x+2)}-\frac{2}{(x+2)(x+1)}\]

myininaya (myininaya):

now look at the bottoms we want the bottoms to be the same so what do we need to multiply the first bottom by

myininaya (myininaya):

what happens if i multiply the first bottom by x+1 then i have to multiply the first top by x+1 \[\frac{x*(x+1)}{(x+3)(x+2)*(x+1)}-\frac{2}{(x+2)(x+1)}\] now what do i need to multiply the second bottom by?

myininaya (myininaya):

hello?

OpenStudy (anonymous):

2

myininaya (myininaya):

we want the bottoms to be the same what do we need to multiply the second bottom by for this to happen?

OpenStudy (anonymous):

is the answer x-2 (x+3) (x+2)

OpenStudy (anonymous):

hello

myininaya (myininaya):

you never answered my question

OpenStudy (anonymous):

by

OpenStudy (anonymous):

by 1x

myininaya (myininaya):

do you see the bottoms? how can we make them the same? what factor is the second fraction missing on the bottom that will make them the same?

OpenStudy (anonymous):

(x+3)

myininaya (myininaya):

thats right so if we multiply the second fraction by (x+3)/(x+3) we will have common denominators so we can combine the fractions like \[\frac{x*(x+1)}{(x+3)(x+2)(x+1)}-\frac{2*(x+3)}{(x+1)(x+2)*(x+3)}=\frac{x(x+1)-2(x+3)}{(x+1)(x+2)(x+3)}\]

myininaya (myininaya):

\[\frac{x^2+x-2x-6}{(x+1)(x+2)(x+3)}\] what do we get on top after combining like terms?

OpenStudy (anonymous):

x^2times -6

myininaya (myininaya):

x^2+x-2x-6 does not equal x^2(-6)

myininaya (myininaya):

x-2x=-x so we have on top x^2-x-6

myininaya (myininaya):

is x^2-x-6 factorable?

OpenStudy (anonymous):

yes

myininaya (myininaya):

what does x^2-x-6 factor to be?

OpenStudy (anonymous):

(x-2)(x+3)

myininaya (myininaya):

close... it should be (x-3)(x+2)

myininaya (myininaya):

so we have \[\frac{(x-3)(x+2)}{(x+1)(x+2)(x+3)}\] does anything cancel?

OpenStudy (anonymous):

x+2

myininaya (myininaya):

right so we have the answer to be \[\frac{x-3}{(x+1)(x+3)}\]

myininaya (myininaya):

got it?

OpenStudy (anonymous):

yea thanks

myininaya (myininaya):

:)